OpenStudy (unklerhaukus):

Prove: \[\int_{-∞}^∞ e^{-\lambda x^2}\text{d}x =\sqrt{ \frac{π}{\lambda}}\]

OpenStudy (turingtest):

what kind of analysis are you doing here?

OpenStudy (turingtest):

I mean, is this from some section where you are doing integrals like this in polar coordinates? 'cause I don't know how to do that...

OpenStudy (unklerhaukus):

hmm well it is a Gaussian, this is just one step in a question i am trying to solve

OpenStudy (unklerhaukus):

This question, i have all the other working i just cannot see how these two terms are equivalent

OpenStudy (turingtest):

Try the new-ish tag feature: @Zarkon @across @JamesJ darn, none of them are online, but they are your best bet for this type of thing you should address your questions at them more often, due to the caliber of you problems

OpenStudy (anonymous):

this is a pain, because you will not find an anti derivative. you have to go into another dimension if i recall. proof is standard, so a google search will probably do it

OpenStudy (unklerhaukus):

i dont know what i would enter into google

OpenStudy (zarkon):

write as a double integral and switch to polar coordinates

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Gaussian_integral

OpenStudy (turingtest):

@Zarkon I knew it! gotta learn how to do those kinds of integrals

OpenStudy (unklerhaukus):

so first step is to recognise it as a even function , then to make in terms of the gamma function, @Zarkon, i will try that also

OpenStudy (unklerhaukus):

\[\int_{-∞}^∞e^{-\lambda x^2}\text{d}x\]\[=2\int_{0}^∞ e^{-\lambda x^2} \text{d} x\] let \[\qquad \lambda x^2=t\]\[\qquad x=\sqrt{t/\lambda} \]\[\qquad \text{d}x=\frac{t^{-1/2}}{2\sqrt\lambda}\text{d}t\]\[={1 \over \sqrt{\lambda}}\int_{0}^∞ t^{-1/2}e^{-t} \text{d} t\]\[= \frac {1} { \sqrt{ \lambda} } \Gamma(1/2) \]\[=\sqrt{\frac {π}{\lambda}}\]